OpenStudy (anonymous):

How do I write this as an equation? P varies directly as q and inversely as the square root of R

OpenStudy (anonymous):

pVR/ q

OpenStudy (ash2326):

P varied directly with q P proportional to q P varies inversely to square root of R P proportional to \(\frac{1}{\sqrt R}\) combining the two so P is proportional to\(\frac{q}{\sqrt R}\) So \[P= k\frac{q}{\sqrt R}\] k is a constant of proportionality